OpenStudy (anonymous):

sqrt{90} + sqrt{250} How would you solve this?

OpenStudy (anonymous):

sqrt{90} + sqrt{250}

OpenStudy (unklerhaukus):

\[\sqrt{90}+\sqrt{250}\] \[=\sqrt{9\times10}+\sqrt{25\times10}\] \[=\sqrt{9}\times\sqrt{10}+\sqrt{25}\times\sqrt{10}\]

OpenStudy (unklerhaukus):

\[=(\sqrt9+\sqrt{25})\sqrt{10}\]

OpenStudy (anonymous):

you have to simplify further

OpenStudy (unklerhaukus):

you can simplify the square roots in the brackets

OpenStudy (anonymous):

Some how the correct answer is \[8\sqrt{10}\]

OpenStudy (unklerhaukus):

\[3\times3=9\] so \(3=\sqrt9\)

OpenStudy (unklerhaukus):

\[5\times5=?\]

OpenStudy (anonymous):

25

OpenStudy (anonymous):

so do they expect you to add 5 +3 = 8 if so whats the formula they used behind that problem?

OpenStudy (unklerhaukus):

\[=(\sqrt{3\times3}+\sqrt{5\times5})\sqrt{10}\] \[=(\sqrt{3^2}+\sqrt{5^2})\sqrt{10}\] \[=(3+5)\sqrt{10}\]

OpenStudy (anonymous):

ty